40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Mon Nov 25 21:46:56 2019, Koji, Summary, PD QE, System Diagram cryo_pd_test.pdfInAsSb_PD_mount_short.PDFPD_pin.pdf
    Reply  Wed Nov 27 20:38:57 2019, Koji, Summary, PD QE, System Diagram P_20191127_203612_vHDR_On.jpg
       Reply  Sun Dec 1 01:32:19 2019, Koji, Summary, PD QE, PD TEC cooling test 20191129191814_IMG_9146.JPGcooling_dark_current.pdf
          Reply  Tue Dec 3 15:32:39 2019, Koji, Summary, PD QE, PD TEC cooling test 
Message ID: 2464     Entry time: Sun Dec 1 01:32:19 2019     In reply to: 2463     Reply to this: 2466
Author: Koji 
Type: Summary 
Category: PD QE 
Subject: PD TEC cooling test 

While I'm still waiting for the proper connector for the vacuum feedthru of the IRLabs cryostat, I have connected to the Dsub9/15 split cable to another Dsub9 connector so that I can test the cooling of the InAsSb sensor in air. Also, the 2004nm laser, a fiber-coupled faraday isolator, and 90:10 beam splitter was moved to the cryostat table and fixed on a black al breadboard. [Attachment 1]

The InAsSb TEC was controlled by the TEC controller of ITC-50. I didn't change the PID parameters of the controller but the temperature nicely setteled to the setpoint. The sensor has a 2.2kOhm thermister. And the max current for the TEC was unknown. The TEC driver had the current limiter of 0.3A and it was not changed for now. With this current limit, the thermistor resistance of 10Kohm was realized. This corresponds to the temperature of about -20degC. According to the data sheet given by Alex, the resistance/temperature conversion is given by the formula

1/T = 7.755e-4 + 3.425e-4*log(R)+1.611e-13*log(R)^3

Questions:

  • What is the max current for the TEC?
  • What is the calibration formula of the thermister (TB04-222) at cryogenic temperature?
    -> Thermistor datasheet link (pdf)

To satisfy the curiosity, the dark current of a (500um)^2 element was measured between -250K and -300K. At -254K, the dark current went down to the level of 40uA (1/15 of the one at the room temp). For the measurement, the bias voltage was set to be 0.5 and 0.6V. However, it was dependent on the diode current. (Probably the bias circuit has the output impedance). This should be replaced by something else.

To Do

  • 2um laser beam setup (w=100um beam)
  • Bias circuit
  • Quick check of the QE and dark noise at -20degC

 

Attachment 1: 20191129191814_IMG_9146.JPG  4.807 MB  | Hide | Hide all
20191129191814_IMG_9146.JPG
Attachment 2: cooling_dark_current.pdf  93 kB  | Hide | Hide all
cooling_dark_current.pdf
ELOG V3.1.3-